OpenStudy (anonymous):

How can you simplify (A. power of 2) power of 6?

OpenStudy (anonymous):

you mean how can you simplify :\[(A^2)^6\]?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

A^12

OpenStudy (anonymous):

\[(A^2)^6 = A^{(2 x 6) }= A^{12}\]

OpenStudy (anonymous):

is it clearer now? ^_^

OpenStudy (angela210793):

A^12

OpenStudy (anonymous):

Rule:\[a^b \times a^c = a^{b+c}\]\[(a)^b(a)^c = a^{bc}\] do you understand the difference now? :)

OpenStudy (anonymous):

yes.(: Thanks.

OpenStudy (anonymous):

np dear ^_^ cheer up :)

OpenStudy (anonymous):

haha What?(:

OpenStudy (anonymous):

smile :)<-- instead of :(<--

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Yeahh.(: